A tennis ball is dropped from 1.21 m above the ground. It rebounds to a height of 1.05 m. With what velocity does it hit the ground? The acceleration of gravity is 9.8 m/s 2 . (Let down be negative.) Answer in units of m/s.

With what velocity does it leave the ground? Answer in units of m/s
1
Expert's answer
2020-09-22T15:36:27-0400

solution

given data

initial velocity when dropped(u)=0 m/s

distance traveled(s)=1.21 m/s

distance traveled after leave the ground(s1)=1.05 ms


according to third law of motion


"v^2=u^2+2gs"


initial velocity u=0


"v^2=2gs\space or\space v=\sqrt{2gs}"


by putting the value of s


"v_{hitting}=\sqrt{2\times9.8\times1.21}=4.86m/s"


after leaving ground at 1.21 m final velocity will be zero

so velocity just after left the ground will be


"v_{2}=\sqrt{2\times9.8\times1.05}=4.53m/s"


so velocity of hitting the ground is 4.86 m/s and after hitting is 4.53 m/s .